OpenStudy (brooke..help00):

@stickos25 @otherworldly The perimeter of a rectangle is 64 units. Can the length x of the rectangle can be 20 units when its width y is 11 units? No, the rectangle cannot have x = 20 and y = 11 because x + y ≠ 64 No, the rectangle cannot have x = 20 and y = 11 because x + y ≠ 32 Yes, the rectangle can have x = 20 and y = 11 because x + y is less than 64 Yes, the rectangle can have x = 20 and y = 11 because x + y is less than 32

jhonyy9 (jhonyy9):

perimeter = 64 P = 2 length +2 width 64 = 2*20 +2*11 64 = 40 +22 64 = 62 so how you think it can being the length 20 units ?

OpenStudy (otherworldly):

what do u think it is

OpenStudy (brooke..help00):

Is it C?

OpenStudy (brooke..help00):

@OtherWorldly

OpenStudy (otherworldly):

i don't know about this one

OpenStudy (stickos25):

It's A